OpenStudy (anonymous):

A segment with endpoints C (3, 4) and D (11, 3) is divided by a point E such that CE and DE form a 3:5 ratio. Find E. Select one: a. (6, 3.625) b. (5.825, 3.7) c. (7.5, 3.4) d. (7, 3.5)

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

Is c=(7.8,3.4)?

OpenStudy (anonymous):

c= 7.5,8.4

OpenStudy (anonymous):

Let M=(3,3), The triangle CMD is a right triangle on M. Let E be the point on CD that divides CD in the given proportion 3:5 Draw a line from E perpendicular the the x-axis. The line intersect MD on H. H has to divide MD in the same proportion 3:5 (similarity) Hence MH=3 and HD =5 Hence H=(6,3) and therefore E=(6,something) Your answer is a

OpenStudy (anonymous):

To find this something, you find the equation of the line CD and you get \[ y= \frac{35-x}{8}\\ y_E= \frac{35-6}{8}=\frac {29}8 = 3.625 \]